Of every 4000 bolts produced in a factory , 800 ofbthem are faulty calculate the probability of a bolt chosen at random not being faulty
1
Expert's answer
2013-02-01T10:00:02-0500
As there are 800 faulty instruments and 4000 of allinstruments, so there are 800 prime events that promotes to choose faulty
instruments, and there 4000 prime events at all. So the probability is 800/4000
= 1/5 to choose faulty instrument.

And probability to choose not faulty is 1 - 1/5 = 4/5
Answer: 4/5